# Revisiting institutional punishment in the $N$-person prisoner’s dilemma

### v2

@@ IV Finite population simulations
Figure 7: Frequencies of cooperators $x$, defectors $y$ and punishers $z$ as a function of time for the initial conditions $z(0)=0.25$ and $x(0)=0.1$ (left panel), $x(0)=0.11052$ (middle panel) and $x(0)=0.3$ ( right panel). The thin curves are the four runs of the stochastic simulation algorithm for $M=10^{4}$ and the thick curves are the numerical solutions of the replicator equations. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.2$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
@@
Figure 8: Probability of fixation of cooperators $\Pi_{C}$ as a function of their initial frequency (left panel) for populations of sizes (from top to bottom) $M=100,500,1000,5000,10000$ and $20000$. The right panel shows $\Pi_{C}$ as a function of the scaled variable $x(0)M^{0.15}$ for $M=5000,10000$ and $20000$. The initial frequency of punishers is set to $z(0)=0.5$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
To quantify the advantage of the cooperators for finite populations, we consider the parameter setting of the simplex plot in the left panel of Fig. 5, where any orbit starting from an initial condition inside the simplex leads to the all-defectors fixed point. The probabilities of fixation $\Pi_{C}$, $\Pi_{D}$, and $\Pi_{P}$ are approximated by the fractions of the runs that lead to the fixation of cooperators, defectors and punishers, respectively. In Fig.  8 we fix $z(0)=0.5$ and vary $x(0)$ in the interval $(0,0.5)$ for several population sizes. The total number of independent runs is $10^{4}$. Since we never observed the fixation of the punishers (i.e., $\Pi_{P}=0$), it is sufficient to show $\Pi_{C}$ in this figure. The results show that fixation of cooperators is very likely for small populations. Even for large populations, the cooperators are very likely to end up as winners, provided, somewhat counterintuitively, that their initial frequency is sufficiently low. This is a consequence of the peculiarity of the trajectories shown in the simplex plot in the left panel of Fig. 5. As $M$ increases, the range of initial conditions $x(0)$ for which $\Pi_{C}$ is not zero shrinks to zero, as expected from the analysis of the replicator equations. More specifically, $x(0)$ decreases with $M^{-0.15}$, as shown in the right panel of Fig. 8. This result implies that, for example, for $x(0)=0.2M^{-0.15}$ and $z(0)=0.5$, the cooperators are virtually certain to fixate regardless of the size of the population. Another important quantity to characterize the stochastic dynamics is the mean fixation time $T_{f}$, i.e., the mean time for the dynamics to reach the all-cooperators or the all-defectors absorbing configurations, which is shown in Fig. 9. For small $x(0)$ and large $M$, the cooperators fixate in practically all runs (see Fig. 8) and the fixation time decreases with increasing $x(0)$, but undergoes an abrupt increase when it is the turn of the defectors to fixate with certainty. In this regime, $T_{f}$ decreases slowly as $x(0)$ increases, which implies that $y(0)$ decreases. In the region of interest, i.e., for $\Pi_{C}\approx 1$ and large $M$, the scaling assumption
@@
with $a=0.15$ and $b=0.1$ fits the data very well, as shown in the right panel of Fig. 9. Here $f(u)$ is a scaling function that satisfies $\lim_{u\to 0}f(u)\to\infty$. The low values of the exponents $a$ and $b$ in this scaling relation indicate the small quantitative effect of the population size on $\Pi_{C}$ and $T_{f}$.
Figure 9: Mean time for fixation of cooperators or defectors $T_{f}$ as a function of the initial frequency of cooperators (left panel) for populations of sizes (from bottom to top) $M=100,500,1000,5000,10000$ and $20000$. The right panel shows the suitability of the scaling assumption (30) for $M=5000,10000$ and $20000$. The initial frequency of punishers is set to $z(0)=0.5$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
To understand how the fixation probability of the cooperators is affected by the choice of the initial frequency of punishers, in Fig. 10 we set $x(0)=0.2M^{-0.15}$ and vary $z(0)$ for different population sizes $M$. As shown before, this scaling guarantees the fixation of the cooperators for $z(0)=0.5$ in the limit of large $M$. The results indicate that there is a threshold for $z(0)$ below which the cooperators have no chance against the defectors. At this threshold, the mean fixation time is maximal. The finite size effects on these properly scaled measures are negligible, except very close to the threshold.
Figure 10: Probability of fixation of cooperators $\Pi_{C}$ (left panel) and scaled mean time for fixation $T_{f}/M^{0.1}$ (right panel) as a function of initial frequency of punishers $z(0)$ for $M=5000,10000$ and $20000$. The initial frequency of cooperators is set to $x(0)=0.2M^{-0.15}$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
@@ V Discussion
In contrast to most public goods game studies, which, following the seminal experimental studies in behavioral economics Fehr_2000 ; Fehr_2002 , consider small groups of players, we focus here on large groups (i.e., communities), since the setting up and maintenance of a sanctioning institution requires a large number of committed individuals. Peer punishment (i.e., vigilantism) is unlikely in communities not only because of the high individual cost of exposure to retaliation, but also because of a well-known phenomenon in social psychology known as the diffusion of responsibility effect: people often fail to cooperate or punish offenders because they believe that others will or should take responsibility for doing so Darley_1968 .
@@
In the community context, we note the striking similarity between the replicator equation approach to public goods games and Wilson’s model of natural selection of populations and communities Wilson_1980 , which focuses on individual traits that affect the fitness of other members of the population, such as behaviors that alter the environment (e.g., pollution and resource depletion). More pointedly, in Wilson’s model the fitness of individuals is determined locally, taking into account interactions within their trait groups, but their chances of reproduction are dictated by competition in the population at large. This is virtually identical to the imitation dynamics of the $N$-person games: a player’s payoff is determined by the composition of the group of $N$ players to which she belongs, but the decision to change her strategy or not is determined by comparison with the payoff of another player randomly selected from the population, as described in Section II. Another point of similarity between the more recent interdisciplinary game theoretic approach to social dilemmas and the traditional biological approach to the evolution of altruism that is worth noting is that the payoffs of cooperators and defectors in the others-only scenario in the absence of punishment (i. e, setting $\beta=0$ in Eqs. (1) and (2)) are identical to the fitness used in Hamilton’s classic work on the innate social behavior of humans Hamilton_1975 , a point overlooked in the recent literature Hannelore_2010 .
In addition to the emphasis on communities, we consider some realistic improvements to the standard formulation of the $N$-person prisoner’s dilemma with punishment, namely the elimination of overpunishment (i.e., the penalties for the different types of offenses are fixed), the high cost of sanctioning institutions (hence a minimum number $K_{m}=\zeta N$ of punishers to set them up), which must be maintained regardless of the presence or absence of offenders, and the division of the maintenance costs among the punishers.
@@
As expected, the replicator equations which describe the imitation dynamics in the infinite population limit (i.e., $M\to\infty$) predict that the all-defectors solution is always stable and that the all-cooperators solution is always unstable. The all-punishers solution is stable if the per capita cost to punishers of contributing to the sanctioning institution is less than the penalty to cooperators, and this cost plus the cost of contributing to public goods is less than the penalty to defectors. In this case, there is a lower threshold $z^{*}$ for the frequency of punishers in the population at large that guarantees the maintenance of the sanctioning institution and, consequently, the disappearance of both types of offenders, i.e., cooperators and defectors. This threshold has a non-trivial dependence on the community size $N$ (see Figs. 1 and 2): $z^{*}$ increases with increasing $N$ when the per capita contribution to the sanctioning institution $\gamma$ is small, and decreases when it is large. For infinitely large communities (i.e., $N\to\infty$), we find $z^{*}\geq\zeta$, as expected.
@@
The imitation stochastic dynamics Fontanari_2024b ; Traulsen_2005 ; Sandholm_2010 for finite populations produces some unexpected results in the case where the replicator equations predict that cooperators win over punishers, who in turn win over defectors if their initial frequency is sufficiently high (see left panel of Fig 5). We find that demographic noise helps to fixate cooperators even for infinitely large populations, provided, somewhat counterintuitively, that the initial fraction of cooperators vanishes with some small power of the population size. Of course, since a population of cooperators can be invaded by a single defector regardless of its size, this is not a robust scenario for maintaining cooperation.
@@
Our approach follows the Hobbesian view that, human beings being the selfish creatures that they are, life would be "solitary, poor, nasty, brutish, and short", were it not for the existence of repressive mechanisms Hobbes_1651 . From a game theory perspective, the main difficulty in establishing this scenario, which seems to be the current state of affairs, is the need for a minimum number of individuals willing to contribute to the establishment of a sanctioning institution. Possible evolutionary trajectories leading to the current repressive dominant system include temporary ritual ceremonies ubiquitous in early human societies, as well as the tendency of human groups to differentiate themselves from neighbors Graeber_2021 . It would be most instructive to incorporate these elements into a game-theoretic framework.
Acknowledgements.

### v3

@@ 4 Finite population simulations
Figure 7: Frequencies of cooperators $x$, defectors $y$ and punishers $z$ as a function of time for the initial conditions $z(0)=0.25$ and $x(0)=0.1$ (left panel), $x(0)=0.11052$ (middle panel) and $x(0)=0.3$ ( right panel). The thin curves are the four runs of the stochastic simulation algorithm for $M=10^{4}$ and the thick curves are the numerical solutions of the replicator equations. The color convention is the same for the three panels. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.2$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
@@
Figure 8: Probability of fixation of cooperators $\Pi_{C}$ as a function of their initial frequency (left panel) for populations of sizes (from top to bottom) $M=100,500,1000,5000,10000$ and $20000$. The right panel shows $\Pi_{C}$ as a function of the scaled variable $x(0)M^{0.15}$. The color convention is the same for both panels. The initial frequency of punishers is set to $z(0)=0.5$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$.
To quantify the advantage of the cooperators for finite populations, we consider the parameter setting of the simplex plot in the left panel of Fig. 5, where any orbit starting from an initial condition inside the simplex leads to the all-defectors fixed point. The probabilities of fixation $\Pi_{C}$, $\Pi_{D}$, and $\Pi_{P}$ are approximated by the fractions of the runs that lead to the fixation of cooperators, defectors and punishers, respectively. In Fig.  8 we fix $z(0)=0.5$ and vary $x(0)$ in the interval $(0,0.5)$ for several population sizes. The total number of independent runs is $10^{4}$. Since we never observed the fixation of the punishers (i.e., $\Pi_{P}=0$), it is sufficient to show $\Pi_{C}$ in this figure. The results show that fixation of cooperators is very likely for small populations. Even for large populations, the cooperators are very likely to end up as winners, provided, somewhat counterintuitively, that their initial frequency is sufficiently low. This is a consequence of the peculiarity of the trajectories shown in the simplex plot in the left panel of Fig. 5. As $M$ increases, the range of initial conditions $x(0)$ for which $\Pi_{C}$ is not zero shrinks to zero, as expected from the analysis of the replicator equations. More specifically, $x(0)$ decreases with $M^{-0.15}$, as shown in the right panel of Fig. 8. This result implies that, for example, for $x(0)=0.2M^{-0.15}$ and $z(0)=0.5$, the cooperators are virtually certain to fixate regardless of the size of the population.
Another important quantity to characterize the stochastic dynamics is the mean fixation time $T_{f}$, i.e., the mean time for the dynamics to reach the all-cooperators or the all-defectors absorbing configurations, which is shown in Fig. 9. We recall that $T_{f}$ is dimensionless and is measured in Monte Carlo steps. For small $x(0)$ and large $M$, the cooperators fixate in practically all runs (see Fig. 8) and the fixation time decreases with increasing $x(0)$, but undergoes an abrupt increase when it is the turn of the defectors to fixate with certainty. In this regime, $T_{f}$ decreases slowly as $x(0)$ increases, which implies that $y(0)$ decreases. In the region of interest, i.e., for $\Pi_{C}\approx 1$ and large $M$, the scaling assumption
@@
with $a=0.15$ and $b=0.1$ fits the data well, as shown in the right panel of Fig. 9. Here $f(u)$ is a scaling function that satisfies $\lim_{u\to 0}f(u)\to\infty$. Note that the scaling form is obeyed only for very large $M$ due to the low values of the exponents $a$ and $b$.
Figure 9: Mean time for fixation of cooperators or defectors $T_{f}$ as a function of the initial frequency of cooperators (left panel) for populations of sizes $M=100,500,1000,5000,10000$ and $20000$. The right panel shows the suitability of the scaling assumption (31). The color convention is the same for both panels. The initial frequency of punishers is set to $z(0)=0.5$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$. $T_{f}$ is dimensionless and is measured in Monte Carlo steps.
To understand how the fixation probability of the cooperators is affected by the choice of the initial frequency of punishers, in Fig. 10 we set $x(0)=0.2M^{-0.15}$ and vary $z(0)$ for different population sizes $M$. As shown before, this scaling guarantees the fixation of the cooperators for $z(0)=0.5$ in the limit of large $M$. The results indicate that there is a threshold for $z(0)$ below which the cooperators have no chance against the defectors. At this threshold, the mean fixation time is maximal.
Figure 10: Probability of fixation of cooperators $\Pi_{C}$ (left panel) and scaled mean time for fixation $T_{f}/M^{0.1}$ (right panel) as a function of initial frequency of punishers $z(0)$ for $M=5000,10000$ and $20000$. The color convention is the same for both panels. The initial frequency of cooperators is set to $x(0)=0.2M^{-0.15}$. The parameters are $N=50$, $K_{m}=10$, $\alpha=0.05$, $\gamma=1.5$, $\beta=3$, $r=2$ and $c=1$. $T_{f}$ is dimensionless and is measured in Monte Carlo steps.
@@ 5 Discussion
In contrast to most public goods game studies, which, following the seminal experimental studies in behavioral economics [49, 50], consider small groups of players, we focus here on large groups (i.e., communities), since the setting up and maintenance of a sanctioning institution requires a large number of committed individuals. Peer punishment (i.e., vigilantism) is unlikely in communities not only because of the high individual cost of exposure to retaliation, but also because of a well-known phenomenon in social psychology known as the diffusion of responsibility effect: people often fail to cooperate or punish offenders because they believe that others will or should take responsibility for doing so [51].
@@
In the community context, we note the striking similarity between the replicator equation approach to public goods games and Wilson’s model of natural selection of populations and communities [52], which focuses on individual traits that affect the fitness of other members of the population, such as behaviors that alter the environment (e.g., pollution and resource depletion). More pointedly, in Wilson’s model the fitness of individuals is determined locally, taking into account interactions within their trait groups, but their chances of reproduction are dictated by competition in the population at large. This is virtually identical to the imitation dynamics of the $N$-person games: a player’s payoff is determined by the composition of the group of $N$ players to which she belongs, but the decision to change her strategy or not is determined by comparison with the payoff of another player randomly selected from the population, as described in Section 2. Another point of similarity between the more recent interdisciplinary game theoretic approach to social dilemmas and the traditional biological approach to the evolution of altruism that is worth noting is that the payoffs of cooperators and defectors in the others-only scenario in the absence of punishment (i. e, setting $\beta=0$ in Eqs. (1) and (2)) are identical to the fitness used in Hamilton’s classic work on the innate social behavior of humans [7], a point overlooked in the recent literature [17].
The study of large groups in finite as well as infinite populations is done in a variant of the standard formulation of the $N$-person prisoner’s dilemma with pool punishment, where there is no overpunishment, i.e, the punishments for the different types of offenses are fixed, and a minimum number $K_{m}=\zeta N$ of punishers is required to form the sanctioning institution, which must be maintained regardless of the presence or absence of offenders. This constant maintenance cost is actually what distinguishes institutional punishment from peer punishment. The burden on punishers is reduced by assuming that the cost of maintenance is shared among punishers once their number exceeds $K_{m}$. All of these features have already been considered in the literature in a variety of different contexts (see, e.g., [25, 26, 27, 28, 37]), but here we have integrated them to offer a more complete model of institutional punishment.
@@
As expected, the replicator equations which describe the imitation dynamics in the infinite population limit (i.e., $M\to\infty$) predict that the all-defectors solution is always stable and that the all-cooperators solution is always unstable. The all-punishers solution is stable if the per capita cost to punishers of contributing to the sanctioning institution is less than the penalty to cooperators, and this cost plus the cost of contributing to public goods is less than the penalty to defectors. In this case, there is a lower threshold $z^{*}$ for the frequency of punishers in the population at large that guarantees the maintenance of the sanctioning institution and, consequently, the disappearance of both types of offenders, i.e., cooperators and defectors. This threshold has a non-trivial dependence on the community size $N$ (see Figs. 1 and 2): $z^{*}$ increases with increasing $N$ when the per capita contribution to the sanctioning institution $\gamma$ is small, and decreases when it is large. Sharing the cost of maintaining the sanctioning institution among the punishers once their number exceeds $K_{m}$ has some notable quantitative effects. To appreciate this, note that the scenario where there is no cost sharing and the punishers’ contribution to the sanctioning institution is $\gamma$, regardless of their number, is obtained by setting $\zeta=1$ in Eqs. (23) and (24), which determine the eigenvalues of the community matrix for the all-punishers solution. Thus, sharing costs (i.e., $\zeta<1$) increases the region of stability of the all-punishers fixed point in the parameter space. Similarly, Eqs. (25) and (26) show that cost sharing also increases the size of the domain of attraction of the all-punishers fixed point in the case of bistability.
We note that the scenario of bistability between punishers and defectors, i.e., the dependence of the equilibrium solution on the initial frequencies of the strategies, is a typical result of well-mixed populations, which is usually not observed in spatial public goods games, where players are fixed at the sites of a lattice and interactions are only possible between neighbors (see, e.g., [19, 20, 21, 22]). In fact, spatial evolutionary games can be quite different from their well-mixed (mean-field) counterparts. For example, in the spatial 2-person prisoner’s dilemma, by randomizing the location of the players, there is a high probability that two (or more) cooperators will be neighbors, resulting in changing spatial patterns where both cooperators and defectors persist indefinitely [53], whereas the well-mixed population approach predicts the rapid extinction of cooperators. Also, the random assignment of mutually dependent players in the same spatial neighborhood allows full invasion of resident populations when mean-field theory predicts bistability [54]. However, in the institutional punishment scenario considered in our paper, where large group sizes $N$ and a minimum number of punishers $K_{m}=\zeta N$ are required to form the punishing institution, spatial localization is unlikely to solve the problem of the dependence on initial frequencies leading to bistability in a well-mixed population: we need to have at least $K_{m}$ punishers in the initial population setting, and the probability that they will all be assigned to sites in the same extended neighborhood is vanishingly small.
@@
The imitation stochastic dynamics [30, 31, 32] for finite populations produces some unexpected results in the case where the replicator equations predict that cooperators win over punishers, who in turn win over defectors if their initial frequency is sufficiently high (see left panel of Fig. 5). We find that demographic noise helps to fixate cooperators even for infinitely large populations, provided, somewhat counterintuitively, that the initial fraction of cooperators vanishes with some small power of the population size. Of course, since a population of cooperators can be invaded by a single defector regardless of its size, this is not a robust scenario for maintaining cooperation.
@@
Our approach follows the Hobbesian view that, human beings being the selfish creatures that they are, life would be ”solitary, poor, nasty, brutish, and short”, were it not for the existence of repressive mechanisms [55]. From a game theory perspective, the main difficulty in establishing this scenario, which seems to be the current state of affairs, is the need for a minimum number of individuals willing to contribute to the establishment of a sanctioning institution. We mention, however, that there are alternative scenarios that actually seem to describe early human societies more accurately, where cooperation is the norm and sanctioning institutions are essentially nonexistent [56].
Acknowledgments
@@
Declaration of interest:
None.
Appendix A
Here we show how the sums in Eqs. (6), (7), and (8) can be explicitly evaluated in the limit of large group size $N$, and hence large $K_{m}=\zeta N$, to derive the discontinuous expected payoffs given in Eqs. (10), (11), and (12).
The difficulty to evaluate the expected payoff of a cooperator $\pi^{C}$ given in Eq. (6) is to calculate the sum
$\Sigma_{C}=\sum_{K=0}^{K_{m}-1}\binom{N-1}{K}z^{K}(1-z)^{N-1-K}$ (A.1)
where the summand is the binomial distribution with parameters $N-1$ and $z$. For large $N$ and $z$ not too small we can replace this distribution by a Gaussian distribution with mean $\mu=Nz$ and variance $\sigma^{2}=Nz(1-z)$,
$\Sigma_{C}$$\approx$ $\int_{0}^{K_{m}}\frac{dk}{\sqrt{2\pi\sigma^{2}}}\exp\left[-\frac{(k-\mu)^{2}}{2\sigma^{2}}\right]$ (A.2)
[…]
$\approx$ $\frac{1}{2}\mbox{erf}\left[\frac{K_{m}-\mu}{\sqrt{2\sigma^{2}}}\right]+\frac{1}{2}\mbox{erf}\left[\frac{\mu}{\sqrt{2\sigma^{2}}}\right].$
Plugging the explicit expressions for $K_{m}$, $\mu$, and $\sigma^{2}$ into this equation yields
$\Sigma_{C}\approx\frac{1}{2}\mbox{erf}\left[\frac{\sqrt{N}(\zeta-z)}{\sqrt{2z(1-z)}}\right]+\frac{1}{2}\mbox{erf}\left[\frac{\sqrt{N}z}{\sqrt{2z(1-z)}}\right],$ (A.3)
so now we can easily take the limit $N\to\infty$. Noticing that $\lim_{x\to\infty}\mbox{erf}(x)\to 1$ and $\lim_{x\to-\infty}\mbox{erf}(x)\to-1$, we get
[…]
Inserting this result into Eq. (6) yields the cooperators’ mean payoff in the limit of infinite group sizes, Eq. (10). Since $\Sigma_{C}$ also appears in Eq. (7), the mean payoff of the defectors in this limit, given in Eq. (11), is also immediately obtained. Finally, the two sums appearing in Eq. (8) are identical to $\Sigma_{C}$ in the limit $N\to\infty$, since $N-1\approx N$ and $K_{m}-2\approx K_{m}$ in this limit, leading directly to the mean payoff of the punishers given in Eq. (12).
Note that the variable $z$ is determined by the replicator equations (16) and (17), and so we have no a priori control over the values it takes in the interval $[0,1]$. In particular, by solving the equations $\pi^{C}(z^{*})=\pi^{D}(z^{*})=\pi^{P}(z^{*})$ for finite $N$ (i.e., numerically evaluating $\Sigma_{C}$ and the other sums appearing in the equations for the strategies’ mean payoffs) we find that the equilibrium coexistence solution $z^{*}$ tends to $\zeta$ as $N$ increases. More precisely, we find that $z^{*}-\zeta\sim 1/\sqrt{N}$. This implies that the argument of the error function in the first term of the rhs of Eq. (A.3) does not diverge in the limit $N\to\infty$. We deal with this problem in the analysis of the coexistence solution by introducing a new variable $\eta=\sqrt{N}(\zeta-z^{*})$ and deriving an equation for it (see Eq. (28)).